Todd operates a business using the cash basis of accounting. At the end of last year, Todd granted permission to switch his sales in account to the accrual method. Last year Todd made 425000 of sales on account and 66000 was uncollected at the end of the year.
What is Todd 481 adjustment for this year.

Riley operates a plumbing business and this year the 3-year old van he used in the business was destroyed in a traffic accident. The van was originally purchased for 23000 and the adjusted basis was 5425 at the end of the accident. Although the van was worth 6300 at the time of accident, insurance only paid Riley 1575 for the loss.
What is the amount of Riley casualty loss deduction

George operates a business that generates revenue of 63 million and allocable taxable income of 1.38 million. Included in the computation of allocable taxable income were deducted expenses of 246500 of business interest and 256,500 of depreciation..
What is the maximum business interest deduction that George will be eligible to claim this year.

Bill operates proprietorship using the cash method of accounting, and this year he received the following :
230 in cash from a customer for service rendered this year.
A promise to pay 174 from a customer for service rendered this year.
Tickets to a football game worth 185 as payment for service performed last year.
A check for 196 for service rendered this year that Bill forgot to cash.
Answer #1

1) Solution: $16,500

Explanation: The positive §481 adjustment for Todd amounts to $66,000 and this year will be allowed 1/4th of the adjustment can be recognized which amounts to $66000/4 = $16,500

2) Solution: $3,850

Explanation: Riley casualty loss deduction = Adjusted basis - Amount paid by insurance = 5,425 - 1,575 = $3,850

3) Solution: $564,900

Working: (1,380,000 + 246,500 + 256,500) * 30% = 564,900

4) Solution: $611

Explanation: 230 + 185 + 196 = $611

Since promise to pay is not property is not income thus excluded